# Unsupervised Domain Adaptation with an Unobservable Source Subpopulation

### v1

@@ 6 Related Work
Firstly, note that the tail bound described in Assumption 1 is intended to hold uniformly for every $n_{1\cdot 0}$ when estimating $\widehat{f}_{k}$ for $k=0,1$. In other words, for each subsample size $n_{1\cdot 0}$, we have a $r_{n_{1\dot{0}}}$ such that the corresponding estimators $\widehat{f}_{k}$ for $k=0,1$ are required to satisfy the stated concentration inequality. This inequality is analogous to Hoeffding’s inequality and provides a non-asymptotic concentration bound on the estimation error. Similar assumptions have also been adopted in recent work (e.g., Maity et al. (2022), Tsybakov & Audibert (2007)).
We next define the Rademacher complexity (Bartlett & Mendelson 2002) that has been frequently used in machine learning literature to establish a generalization bound. Instead of considering the Rademacher complexity on $\mathcal{F}$ we define the class of weighted losses $\mathcal{G}(\ell,\mathcal{F})=[w(x,y)\ell\{g(x),y\}:g\in\mathcal{F}]$ and $n\in N$ we define its Rademacher complexity measure as
@@
for $\beta_{10}\in(0,\varrho)$ in Wellner et al. (2013)(see Corollary 3.2.3) to conclude that $\widehat{\beta}_{10}\rightarrow\beta_{10}$ in probability and hence $\widehat{\beta}_{10}$ is a consistent estimator for $\beta_{10}$.
@@
Next, we use a symmetrization argument (see Wellner et al. (2013), Chapter 2, Lemma 2.3.1) to bound the expectation $E\{F({\bf Z}_{1:n_{1\cdot 0}})\}$ by the Rademacher complexity of the hypothesis class $\mathcal{G}$, i.e.,
@@
Remark 1.
S.3 Synthetic Data Results
For model estimation, we utilize the known data-generating distribution to compute the conditional probability models required by both our proposed estimators and the benchmark method. Specifically, we calculate the five key conditional probabilities needed for implementation: $\xi_{0}({\bf x})$, $\xi({\bf x})$, $\tau_{0}({\bf x})$, $\tau_{1}({\bf x})$, and $\kappa({\bf x})$, which together determine the models $\eta_{0}({\bf x})$, $\eta_{1}({\bf x})$ and $\eta({\bf x})$ for our method, and $\xi_{0}({\bf x})$, $\xi_{1}({\bf x})$ and $\xi({\bf x})$ for the benchmark. Additionally, the parameters $\{\beta_{ya}:y=0,1;a=0,1\}$ and $\{\alpha_{ya}:y=0,1;a=0,1\}$ as known. The classification threshold is set to $0.5$, consistent with the standard Bayesian decision rule—that is, a sample is classified as positive if the predicted probability exceeds $0.5$.
S.4 Additional Benchmark Data Results

### v2

@@ 8 Discussions and Conclusions
We define the Rademacher complexity (Bartlett and Mendelson 2002) that has been frequently used in machine learning literature to establish a generalization bound. Instead of considering the Rademacher complexity on $\mathcal{F}$ we define the class of weighted losses $\mathcal{G}(\ell,\mathcal{F})=[w(x,y)\ell\{g(x),y\}:g\in\mathcal{F}]$ and $n\in N$ we define its Rademacher complexity measure as
@@
for $\beta_{10}\in(0,\varrho)$ in Wellner and others (2013)(see Corollary 3.2.3) to conclude that $\widehat{\beta}_{10}\rightarrow\beta_{10}$ in probability and hence $\widehat{\beta}_{10}$ is a consistent estimator for $\beta_{10}$.
@@
Next, we use a symmetrization argument (see Wellner and others (2013), Chapter 2, Lemma 2.3.1) to bound the expectation $E\{F({\bf Z}_{1:n_{1\cdot 0}})\}$ by the Rademacher complexity of the hypothesis class $\mathcal{G}$, i.e.,
@@
S.3 Additional Benchmark Data Results